OpenStudy (anonymous):

Tonya is twice Kevin's age. In 3 years, Tonya will be 17. How old is Kevin

OpenStudy (anonymous):

2k = t t + 3 = 17

OpenStudy (anonymous):

A.6 B.20 C.7 D.8

OpenStudy (anonymous):

I think its D

OpenStudy (anonymous):

2k + 3 = 17 2k + 3 -3 = 17 -3 2k + 0 = 14 2k = 14 2k / 2 = 14 / 2 k = 7

OpenStudy (anonymous):

Kevin is seven years of age. I concur with aaron.

OpenStudy (anonymous):

Okay thanks c:

OpenStudy (anonymous):

Kevin tiene siete anos. Estoy de acuerdo con aaron.